OpenStudy (anonymous):

How do I find value of x without using quadratic equation for this 4 = x / (1-x) ^2

OpenStudy (anonymous):

I said I was going to go to sleep, but here are the solutions. x = 1/8 (9-sqrt(17)) x = 1/8 (9+sqrt(17)) Also you said you don't want to solve using quadratic, so set it up a different way. Like so. 4=x/(x-1)^2 Or... 64/17(x-9/8)^2=1

OpenStudy (anonymous):

I just noticed, I repeated one, the second one though works.